Milden Company is a distributor who wants to start using a contribution format income statement for planning purposes. The company has analyzed its expenses and developed the following cost formulas: Cost Cost Formula Cost of good sold $33 per unit sold Advertising expense $189,000 per quarter Sales commissions 8% of sales Shipping expense ? Administrative salaries $99,000 per quarter Insurance expense $10,900 per quarter Depreciation expense $69,000 per quarter Because shipping expense is a mixed cost, the company needs to estimate the variable shipping expense per unit sold and the fixed shipping expense per quarter using the following data: Quarter Units Sold Shipping Expense Year 1: First 35,000 $ 179,000 Second 37,000 $ 194,000 Third 42,000 $ 236,000 Fourth 38,000 $ 199,000 Year 2: First 36,000 $ 189,000 Second 39,000 $ 204,000 Third 53,000 $ 251,000 Fourth 50,000 $ 227,000 Required: 1. Using the high-low method, estimate a cost formula for shipping expense in the form Y = a + bX. 2. In the first quarter of Year 3, the company plans to sell 41,000 units at a selling price of $66 per unit. Prepare a contribution format income statement for the quarter.
Answer #1

Answer

  • Working for requirement 1

period

Units

Cost

High Level

Year 2, 3rd Quarter

                          53,000

$                251,000.00

Low Level

Year 1, 1st quarter

                          35,000

$                179,000.00

Difference

                          18,000

$                   72,000.00

  • Requirement 1

Difference in Cost

$                                                             72,000.00

Difference in units

                                                                      18,000

Variable cost per unit

$                                                                       4.00

Working

High Level

Low Level

A

Total Cost

$              251,000.00

$                179,000.00

B

Total Units

53000

35000

C

Variable cost per unit

$                           4.00

$                             4.00

D = B x C

Total Variable cost

$              212,000.00

$                140,000.00

E = A - D

Total Fixed Cost

$                 39,000.00

$                   39,000.00

Hence, Variable shipping expense = $ 4 per unit
Fixed shipping expense = $ 39000 per quarter

Cost Formula for shipping expense: Y = a +bX
Y [total shipping expense] = $ 39,000 + ($4.00 x Units)

  • Requirement 2

Contribution margin format income statement

A = 41000 units x $ 66

Sales Revenue

$2,706,000

Variable costs:

B =41000 units x $33

Cost of Goods Sold

$1,353,000

C = A x 8%

Sales Commission

$216,480

D = 41000 units x $ 4

Shipping Expenses

$164,000

E = B+C+D

Total variable cost

$1,733,480

F = A - E

Contribution margin

$972,520

Fixed Costs:

Advertising expense

$189,000

Shipping Expenses

$39,000

Administrative salaries

$99,000

Insurance expense

$10,900

Depreciation expense

$69,000

G

Total Fixed Cost

$406,900

H = F - G

Net Income

$565,620